Update - We have again been continuing to monitor this closely overnight and have seen no further attacks since 18:20 BST 21/06/21. Traffic patterns on the network appear normal and services should be operational. Any customers still seeing an issues please raise an incident with the Support team to investigate.
Jun 23, 08:14 BST
Update - We have been continuing to monitor this closely overnight and have seen no further attacks since 18:20 BST 21/06/21. Traffic patterns on the network appear normal this morning and services should be operational. Any customers still seeing an issues please raise an incident with the Support team to investigate.
Jun 22, 11:41 BST
Monitoring - Further mitigations were deployed and we saw traffic levels drop to normal levels at 18:20 BST. We are continuing to monitor for further waves of attack on the network.
Jun 21, 22:00 BST
Update - At approx 17:50 we started to experience another wave of attack. We are currently putting in place further mitigations against this traffic.
Jun 21, 18:14 BST
Update - We are continuing to work on this issue. Currently traffic levels are back down to normal levels and we have mitigations currently in place. Any customers still seeing DNS issues from this may want to temporarily try a public DNS server whilst this incident is ongoing.
Jun 21, 15:29 BST
Identified - We are seeing a large increase in DDoS traffic - we are working to mitigate this new wave
Jun 21, 12:00 BST
Update - We are seeing our mitigation working - and traffic levels have reduced significantly. We are continuing to monitor closely and work with our upstream transit providers to filter the traffic.
Jun 21, 09:44 BST
Update - We are seeing another DDOS wave and are blackholing traffic to affected IPs. We will provide another update shortly.
Jun 21, 08:41 BST
Update - We are continuing to monitor for any further issues.
Jun 20, 20:41 BST
Monitoring - We are seeing normal traffic levels continue and services remain stable since our upstream carriers applied the requested filters and blackhole routing.

We are keeping this under monitor
Jun 20, 20:39 BST
Identified - We have been liaising with our upstream carriers and they have applied some filters and blackhole routing to mitigate the attack. We are gradually seeing traffic levels drop and some services restore.
Jun 20, 18:45 BST
Investigating - We are currently experiencing major packet loss and interruptions due to a suspected DDoS attack on the amatis Network.
Jun 20, 17:57 BST
Connectivity Operational
Ethernet Services (Fibre, EFM, GEA) ? Operational
LTE (4G) Services ? Operational
xDSL Services (ADSL/FTTC) ? Operational
Colocation (Network Connectivity) Operational
Cloud Services Operational
Virtual Datacentre (VDC) Operational
Virtual Private Servers (VPS) Operational
High Speed Storage Services ? Operational
Archival Backup Storage Services ? Operational
Core Partial Outage
Core Network Devices and Routing ? Operational
Internet Transit ? Operational
Public Internet Peering ? Partial Outage
Private Internet Peering ? Operational
Infrastructure Operational
Datacentre Issues ? Operational
Power ? Operational
Cooling ? Operational
Security ? Operational
Platform Services Operational
Service Availability Monitoring ? Operational
DNS Services ? Operational
amatis Email Services ? Operational
Customer Portals, Tools and Contact Operational
Customer Graphing Tools ? Operational
amatis Phone System ? Operational
Customer Support Portal ? Operational
amatis Website ? Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Jun 25, 2021

No incidents reported today.

Jun 24, 2021
Resolved - Our LTE provider have confirmed the issue within the cell network is repaired and we are seeing 4G services stable since 23/06/21 15:25. Any customers still experiencing issues with their service please get in touch.
Jun 24, 08:47 BST
Monitoring - Our provider have advised that this fault is now resolved and we are seeing connections back online. The issue appears to be on the cell provider network and we are waiting the root cause.
Jun 23, 16:37 BST
Identified - Our LTE provider have confirmed there is a major service outage on their network, however, no ETR has been provided as yet. We are continuing to push for this.
Jun 23, 13:09 BST
Update - 4G services remain down. Our LTE provider are still investigating the fault
Jun 23, 11:32 BST
Investigating - We are currently investigating a fault with LTE/4G services and have raised the issue to our provider to investigate.
Jun 23, 08:11 BST
Jun 23, 2021

Unresolved incident: INC0035836 - Suspected DDoS.

Jun 22, 2021
Jun 21, 2021
Jun 20, 2021
Jun 19, 2021

No incidents reported.

Jun 18, 2021
Postmortem - Read details
Jun 18, 10:37 BST
Resolved - This incident has been resolved.
Jun 18, 10:34 BST
Monitoring - Services should now be restored.
The issue was a large distributed DOS attack - which has now been mitigated.
A full RFO will be issued in due course.
Jun 15, 17:31 BST
Investigating - Hi,
We are investigating some disruption events across our network. We will post an update here shortly.
Jun 15, 16:01 BST
Jun 17, 2021

No incidents reported.

Jun 16, 2021

No incidents reported.

Jun 15, 2021
Jun 14, 2021

No incidents reported.

Jun 13, 2021

No incidents reported.

Jun 12, 2021

No incidents reported.

Jun 11, 2021
Completed - The router has been successfully replaced.
If any customer connected to Reading South exchange experiences any issues please contact amatis support on 01183219946 or email support@amatisnetworks.com
Jun 11, 21:37 BST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jun 11, 19:32 BST
Scheduled - amatis will be carrying out emergency maintenance to replace a faulty router in the APEX Reading South exchange. Customers that terminate on this device will see a brief loss of service whilst the service is migrated on to the replacement device.
Jun 11, 17:21 BST